# MotionDreamer: Zero-Shot 3D Mesh Animation from Video Diffusion Models

### v1

@@ Conclusion
We presented a novel generative method for zero-shot 3D animation. Despite its limitations stemming from the single-view supervision, we demonstrated that it produces visually preferable motions across diverse unseen 3D shapes. We see our method as a capable tool for analysis of motion spaces in VDMs, and we envision its future application for affordable re-animation of otherwise static 3D assets in large-scale virtual environments.
@@
Our method produces novel poses for 3D objects including human bodies and faces, which are sensitive to misuse. However, we neither focus on this class nor do we strive for realistic appearance modeling. Furthermore, the animations produced by our method are influenced by biases of the backbone VDMs, and we conclude that their research is a matter of high interest to the community.
@@ B.1 Baseline methods
We observe a lack of class-agnostic end-to-end pure motion generators. Therefore, we combine a human-specific motion generator with a general motion transfer method while accepting that the performance of such solution will depend on morphological and semantic proximity of the source and target shape class. To this goal, we first use a pre-trained author’s implementation of the text-conditioned motion diffusion by Tevet et al. [79] to generate a unique 2D skeletal human motion sequence for each example in our study. We adapt the motion text prompts used for our method (Tbl. 2) to the human domain using a template “a person is [ACTION]” e.g., “a horse is walking” $\to$ “a person is walking”. Next, we use a 2D-to-3D human body pose uplifting method adapted from the code of Zuo et al. [101] to obtain sequence of SMPL [27] meshes. Finally, we follow the procedure and code of Liao et al. [84] to retarget the SMPL animations to our target meshes. We aply this step consistently even for the SMPL target mesh. Finally, we render the first 16 frames of the resulting mesh sequences in the same way as for our own method.
@@ B.2 Stimuli
We extract the surface models for SMPL [27], FLAME [28] and SMAL [29] from their official implementations. For SMPL, we opt to lower the arms to 45 degrees from the original T-pose, while we use the default “zero” pose parameters for others. For NJF [26], we use the 4 open assets listed in Tbl. 1.
@@
Tbl. 2 lists VDM prompts used to generate the motion sequences for the stimuli in our study.
@@
Tbl. 3 shows the positive and negative prompts used for Single-view Texturing as an input for the ControlNet diffusion model [89] for each shape in our experiments and the prompt for the Stable Diffusion XL [91] background inpainting.
@@ B.4 Detailed Results
Here, we present a break-down of the results from our user study separately for the human stimuli (Fig. 11), where the human-specific MDM-MT baseline performs well and for the remaining stimuli (Fig. 12), where our class-agnostic method dominates. We also offer a detailed breakdown in Tbl. 4.
@@ C.3 WHAM baseline
To offer a fair comparison, we evaluate four different alignment strategies for WHAM [67], because our method starts with the known pose $\mathbf{p}_{\textrm{init}}$. Results for either strategy can be found in  Tbl. 5. In strategy align, we find the rotation and translation to align the wham output with the ground truth:

### v2

@@ Conclusion
We presented a novel generative method for zero-shot 3D animation. Despite its limitations stemming from the single-view supervision, we demonstrated that it produces visually preferable motions across diverse unseen 3D shapes at computation cost lower than end-to-end 4D generative methods. We see our method as a capable tool for analysis of motion spaces in VDMs, and for affordable re-animation of otherwise static 3D assets in large-scale virtual environments.
@@
Our method produces novel poses for 3D objects including human bodies and faces, but we do not focus on realistic appearance modeling. The biases in backbone VDMs can influence our method and are a priority research interest to the community.
@@
MotionDreamer: Exploring Semantic Video Diffusion features for Zero-Shot 3D Mesh Animation
Lukas Uzolas    Elmar Eisemann    Petr Kellnhofer
Delft University of Technology
The Netherlands
@@ B.1 Baseline methods
We observe a lack of class-agnostic end-to-end pure motion generators. Therefore, we combine a human-specific motion generator with a general motion transfer method while accepting that the performance of such solution will depend on morphological and semantic proximity of the source and target shape class. To this goal, we first use a pre-trained author’s implementation of the text-conditioned motion diffusion by Tevet et al. [80] to generate a unique 2D skeletal human motion sequence for each example in our study. We adapt the motion text prompts used for our method (Tbl. 4) to the human domain using a template “a person is [ACTION]” e.g., “a horse is walking” $\to$ “a person is walking”. Next, we use a 2D-to-3D human body pose uplifting method adapted from the code of Zuo et al. [103] to obtain sequence of SMPL [31] meshes. Finally, we follow the procedure and code of Liao et al. [86] to retarget the SMPL animations to our target meshes. We aply this step consistently even for the SMPL target mesh. Finally, we render the first 16 frames of the resulting mesh sequences in the same way as for our own method.
@@ B.2 Stimuli
We extract the surface models for SMPL [31], FLAME [3] and SMAL [1] from their official implementations. For SMPL, we opt to lower the arms to 45 degrees from the original T-pose, while we use the default “zero” pose parameters for others. For NJF [2], we use the 4 open assets listed in Tbl. 3.
@@
Tbl. 4 lists VDM prompts used to generate the motion sequences for the stimuli in our study.
@@
Tbl. 5 shows the positive and negative prompts used for Single-view Texturing as an input for the ControlNet diffusion model [91] for each shape in our experiments and the prompt for the Stable Diffusion XL [93] background inpainting.
@@ B.4 Detailed Results
Here, we present a break-down of the results from our user study separately for the human stimuli (Fig. 16), where the human-specific MDM-MT baseline performs well and for the remaining stimuli (Fig. 16), where our class-agnostic method dominates. We also offer a detailed breakdown in Tbl. 6.
@@ C.2 Methodology
Note that we utilize MSE loss when using RGB features for optimization, as this results in better performance compared to the cosine distance.
@@ C.3 WHAM baseline
To offer a fair comparison, we evaluate four different alignment strategies for WHAM [68], because our method starts with the known pose $\mathbf{p}_{\textrm{init}}$. Results for either strategy can be found in  Tbl. 7. In strategy align, we find the rotation and translation to align the wham output with the ground truth:
@@ D.1 Effect of Texturing and Background for Different VDMs
D.2 Qualitative Comparison to Consistent4D
We further compare our method to Consistent4D [25]. Since this is a computationally significantly more expensive method (50 minutes in its low VRAM setup versus less than 3 minutes for ours), which ened-to-end produces both the shape and the animation, we consider it a separate category from method which is better suited for quick motion prototyping and iterative animation development. This is why we did not include Consistent4D in our user study and instead provide a general discussion and a qualitative comparison here. A similar method DG4D was included in our study instead.
As seen in Fig. 19 and Fig. 20, our method generally produces more plausible motion given the underlying geometry in a faster manner. The evaluation of Consistent4D take on average 32 minutes per object in its low VRAM setup, while our motion fitting takes on average under 3 minutes on an NVIDIA RTX 3090. Note that we exclude the time it takes to generate the driving videos in both cases. Consistent4D’s slower runtime can be explained by its adoption of SDS which necessitates encoding the rendered RGB image into the latent space repeatedly. In contrast, our method remains in the semantic feature space.
Consistent4D utilizes a K-Plane [104] for their 4D representation which allows for a higher flexibility when modeling geometry distortions produced by the VDM. One such example can be seen in the raptor sequence in Fig. 19. Here, the raptor’s right and left legs rapidly alternate between the foreground and background, creating a sense of motion, but disrupting the raptor’s underlying topology. While the ability to fit such non-physical effects can be beneficial in some scenarios, our explicit representation is more robust to fitting to VDM artifacts and thus reduces the consequent visually implausible transformations.
Furthermore, the K-plane representation entangles shape and motion and hence it cannot be easily integrated into common computer-graphics pipelines. This is in contrast to our method which deforms an explicit canonical mesh, where the time-dependent vertex deformations can be easily exported.
Lastly, Consistent4D (as well as DG4D) adopts an image-to-3D model Zero-1-to-3 [84] for 3D-Uplifting. However, Zero-1-to-3 requires input images without background which contradicts our observations that VDMs benefit from context in the background for better results (see in Appendix D.1). To tackle this, Consistent4D creates masks in an automated fashion for videos with background which can, however, introduce additional errors. In comparison, our method does not rely on any such masks.
Figure 19: Comparing Consistent4D against our method.
Figure 20: Comparing Consistent4D against our method.
Figure 21: Comparison of semantic features against RGB used for pose optimization. Each column shows frames for a single sequence. Note the red boxes, highlighting errors in the the pose optimization when utilizing RGB: 1) In case of SMPL (human), the hand gets stuck in front of the torso, as the RGB features do not distinguish the body from the hand. 2) In case of SMAL (horse), the limbs of the horse assume less realistic articulation with RGB features.
D.3 Comparing Semantic Featuring against RGB for Optimization
Here, we complement our Pose estimation experiment from the main paper and compare the RGB and semantic features end-to-end in our full pipeline We find that modeling temporal deformations with NJF in combination with RGB features results in unstable optimization. Therefore, we only showcase the kinematic models in Fig. 21. Similarly to pose fitting, semantic features lead to superior results.
D.4 Ablation of Number of Vertices
Fig. 22 and  Fig. 23 show additional results when varying the number of vertices in our method with NJF. We find that the output quality degrades gracefully and predictably, when scaling down from 4000 to 500 vertices. Notice that the VDM output slightly differs, due to the variations in the conditioning input image when varying the number of vertices.
Figure 22: Effect of vertex number on our method when adopting NJF for deformations with VC as VDM backbone. Each row shows the output of our method with an increasing number of vertices. Prompt: ’An orc laughing.’
Figure 23: Effect of vertex number on our method when adopting NJF for deformations with DC as VDM backbone. Each row shows the output of our method with an increasing number of vertices. Prompt: ’A truck moving its shovel up and down.’
Figure 24: Results when motion fitting using the SVD semantic features. While our method is well suited to work even with SVD features, SVD videos tend to focus on a global camera rotation rather than actual object motion. Consequently, the fitted object motion is often minimal or uninteresting. As a result we omitted SVD in our further studies in favor of other VDMs.
D.5 Results with Stable Video Diffusion
As reported in Appendix D.1, SVD produces camera motion rather than object motion. For completeness sake, we show that our method produces plausible results in Fig. 24 when fitting to the SVD semantic features.